# Filtering for Creativity: Adaptive Prompting for Multilingual Riddle Generation in LLMs

### v1

@@ Abstract
Multilingual riddle generation challenges large language models (LLMs) to balance cultural fluency with creative abstraction. Standard prompting strategies—zero-shot, few-shot, chain-of-thought—tend to reuse memorized riddles or perform shallow paraphrasing. We introduce Adaptive Originality Filtering (AOF), a prompting framework that filters redundant generations using cosine-based similarity rejection, while enforcing lexical novelty and cross-lingual fidelity. Evaluated across three LLMs and four language pairs, AOF-enhanced GPT-4o achieves 0.177 Self-BLEU and 0.915 Distinct-2 in Japanese, signaling improved lexical diversity and reduced redundancy compared to other prompting methods and language pairs. Our findings show that semantic rejection can guide culturally grounded, creative generation without task-specific fine-tuning.
@@ 1 Introduction
Large Language Models (LLMs) excel across many natural language processing tasks but often falter in creative generation within multilingual settings (Zhang (2025); Ismayilzada et al. (2024)). This limitation is especially apparent in bilingual riddle generation, where outputs are frequently rote and easily found online. To address this, we propose Adaptive Originality Filtering (AOF)—a prompting framework that promotes cultural specificity and lexical novelty without fine-tuning. AOF employs a semantic rejection mechanism to discard formulaic outputs and iteratively prompt the model until it produces a response that is both original and culturally coherent. In contrast to existing strategies like zero-shot, few-shot Brown et al. (2020a), and chain-of-thought prompting Wei et al. (2023), which often yield paraphrased common riddles, AOF actively steers generation toward novelty and cultural fidelity.
Figure 1: This diagram illustrates the end-to-end pipeline for generating and validating riddles using LLMs(GPT-4o, R1, LLaMA). The process begins with enforcing constraints on novelty and structure. Model outputs are then checked for semantic similarity using MiniLM, with a threshold score of << 0.75 determining acceptance. Outputs that fail are regenerated (Retry), while accepted samples undergo final evaluation.
We ask: Can prompting alone enable culturally aware, original bilingual generation in LLMs? To explore this, we apply AOF to three state-of-the-art LLMs (GPT-4o, LLaMA 3.1, DeepSeek) across four language pairs, benchmarking against four standard prompting baselines. We evaluate using metrics for lexical diversity (Distinct-n), redundancy (Self-BLEU), bilingual semantic alignment (Cross-Lingual BERTScore), and syntactic validity. Results show that AOF significantly improves both originality and cross-cultural alignment. By injecting an external filtering loop into prompting, AOF offers a lightweight, scalable path toward more creative and culturally attuned multilingual generation (see Figure 1).
@@ 2 Related Work
Most prior research emphasizes riddle solving rather than generation. Panagiotopoulos et al. (2024) and Heavey et al. (2024) developed multilingual inference models for riddle solving but did not extend to generative tasks. Xu et al. (2022) incorporated cultural embeddings for Chinese riddle comprehension, while Smith et al. (2022) evaluated NLG models on static benchmarks without multilingual or generative scope.
Multilingual representation efforts like MUSE Lample & Conneau (2019), LASER Chen & Avgustinova (2021), and XLM-R Conneau et al. (2019) support cross-lingual transfer but not figurative or culturally grounded generation. Sentence-level encoders Xia et al. (2019); Gritta & Iacobacci (2021) improve alignment but lack capacity for creative language. Dufter (2021) noted that such embeddings often collapse metaphorical nuance essential to cultural generation.
Creative Language Generation
Creative generation research has largely focused on ideation and open-domain tasks Ma et al. (2025). Cox et al. (2021; 2023) used structured prompts for idea diversity, while Huang et al. (2023) improved multilingual output via cross-lingual-thought prompting without filtering. Constraint-based methods like Yang et al. (2022) and Laverghetta Jr et al. (2024) use decoding control or heuristics to boost diversity.
Structural prompting has also gained traction: de Wynter et al. (2023) introduced recursive meta-prompts, Verma et al. (2024) identified abstraction limits in ReAct prompting, and Liu et al. (2024) applied paraphrasing under semantic constraints. Atmakuru et al. (2024) benchmarked narrative creativity under fixed templates, highlighting rejection-based filtering as a viable originality enhancer.
Prompting for Originality and Structure
Standard approaches prioritize structure over creativity. Few-shot prompting Brown et al. (2020a; b) and CoT Wei et al. (2023) improve reasoning but often replicate training artifacts. Iterative methods like Self-Refine Madaan et al. (2023), Tree-of-Thought Yao et al. (2023), and Reflexion Krishna et al. (2023) enhance fluency and factuality through retries but rarely enforce novelty.
Constraint-driven prompting offers stronger generative control. COLD decoding Mou et al. (2022), EditCoT Wang et al. (2024), and Sketch-of-Thought Aytes et al. (2025) use template-based completions, while DeLorean Liu et al. (2023) integrates symbolic constraints. These techniques target structure but overlook cultural fidelity. In contrast, our method enforces novelty during decoding via cosine-based filtering and supports cross-lingual generation of culturally resonant riddles.
@@ 3 Methodology
This section describes the components of our multilingual riddle generation framework, including our originality-enforcing prompting method, a centralized automatic evaluation metric, human annotation protocols, experimental design, and fine-tuning our model.
3.1 Adaptive Originality Filtering (AOF) Prompting
Traditional prompting strategies such as Chain-of-Thought (CoT) and Few-Shot improve logical coherence but frequently fall short in producing structurally novel and semantically diverse riddles. Prior work has shown that LLMs often reproduce internet riddles or paraphrase entries from BiRdQA Zhang & Wan (2022), defaulting to overused scaffolds such as “I have…” and “What am I?”. To address this, we propose Adaptive Originality Filtering (AOF), a prompting method designed to enforce semantic novelty, syntactic variation, and cross-lingual fidelity.
Semantic Filtering.
Each generated riddle $r_{\text{gen}}$ is compared against a reference set $\mathcal{D}=\{r_{i}\}_{i=1}^{N}$ using cosine similarity in embedding space. A candidate is accepted only if:
$\mathcal{S}(r_{\text{gen}},\mathcal{D})<\theta,\text{ with }\theta=0.75.$
The full formulation and reasoning for .75 appears in Appendix J.1.
Rejection Sampling Loop.
AOF employs rejection sampling: if a candidate fails the novelty filter, it is discarded and regenerated. This process repeats until a valid output is produced or a retry cap is reached. See Appendix J.2 for pseudocode.
Prompt Constraints.
Prompts used in AOF enforce specific closure formats, grammatical variation, and discourage common answers. Template rules are detailed in Appendix J.
@@ 3.2 Experimental Setup
We evaluate generation across multiple LLMs and prompting strategies, comparing structural, semantic, and fluency metrics in five languages.
Models Evaluated.
We benchmark GPT-4o, LLaMA 3.1, and DeepSeek Reasoning (R1), all queried under identical settings (temperature 0.7, token limit 3000).
Prompting Strategies.
We compare Zero-Shot, Few-Shot, Chain-of-Thought, Adversarial Wallace et al. (2019); Ribeiro et al. (2018), and AOF prompting. Full details are provided in Appendix K.
Dataset.
BiRdQA Zhang & Wan (2022) provides a bilingual corpus of 15,365 riddles in English and Chinese. We sample balanced subsets and align multilingual answers.
Metrics.
Automatic metrics include Self-BLEU (repetition), Distinct-2 (diversity), BERTScore (alignment). Human validation is included in Appendix A.
3.3 Fine-Tuning of the GPT-4o Model
@@
This fine-tuning aimed to improve GPT-4o-2024-08-06’s performance on riddle comprehension and generation across multiple languages. Riddles involve more than surface-level matching—they require metaphor understanding, logical contradiction, and creative misdirection. Our goal was not only to raise answer accuracy but also to instill structural reasoning abilities.
@@
We framed the task as a supervised multi-class classification problem using the BiRdQA dataset. Riddles were presented in a multiple-choice format, and the model was fine-tuned using cross-entropy loss. For complete details on dataset preprocessing, training setup, and training set expansion, see Appendix I.
@@
Riddles were framed in a four-option multiple-choice format to encourage fine-grained discrimination between plausible distractors. This setup shaped the model’s reasoning strategies and generalization capabilities. A full analysis of framing effects is provided in Appendix I.5.
@@
We evaluated five prompting strategies on the fine-tuned model: Zero-Shot, Few-Shot, Chain-of-Thought (CoT), Adversarial, and Adaptive Originality Filtering (AOF). These were held consistent with the pretrained experiments. See Table 21 for full prompt templates.
@@
We compared the fine-tuned GPT-4o to several pretrained baselines—GPT-4o (pretrained), LLaMA 3.1, and DeepSeek R1—using identical prompts and evaluation metrics. For detailed results and methodological discussion, see Appendix I.4.

### v3

# Adaptive Originality Filtering: Rejection‑Based Prompting and RiddleScore for Culturally Grounded Multilingual Riddle Generation

@@ Abstract
Language models are increasingly tested on multilingual creativity, demanding culturally grounded, abstract generations. Standard prompting methods often produce repetitive or shallow outputs. We introduce Adaptive Originality Filtering (AOF), a prompting strategy that enforces novelty and cultural fidelity via semantic rejection. To assess quality, we propose RiddleScore, a metric combining novelty, diversity, fluency, and answer alignment. AOF improves Distinct-2 (0.915 in Japanese), reduces Self-BLEU (0.177), and raises RiddleScore (up to +57.1% in Arabic). Human evaluations confirm fluency, creativity, and cultural fit gains. However, improvements vary: Arabic shows greater RiddleScore gains than Distinct-2; Japanese sees similar changes. Though focused on riddles, our method may apply to broader creative tasks. Overall, semantic filtering with composite evaluation offers a lightweight path to culturally rich generation—without fine-tuning.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP) across a spectrum of applications, yet their generative abilities in creative, multilingual contexts remain underexplored and underperforming (Zhang and Wan, 2025; Ismayilzada et al., 2024). Tasks like riddle generation pose a unique challenge: success hinges not only on linguistic fluency but also on metaphorical abstraction, cultural resonance, and semantic ambiguity—all of which are frequently underrepresented in LLM training corpora (Sejnowski, 2023; Pawar et al., 2024). As LLMs are increasingly integrated into global educational and creative platforms, their limitations in culturally grounded generation constrain both inclusivity and expressive potential(Bulathwela et al., 2024; Spennemann, 2023).
Figure 1: End-to-end pipeline to produce and verify riddles with LLMs (GPT-4o, R1, LLaMA). Constraints enforce novelty/structure; MiniLM tests semantic similarity with threshold $\leq 0.75$. Failed results are re-generated; accepted ones are subjected to final checking.
Riddles, with their blend of metaphor, misdirection, and context-specific symbolism, provide a compelling benchmark for evaluating multilingual creativity in NLP. However, existing prompting strategies—zero-shot, few-shot, and chain-of-thought—often yield formulaic outputs or mistranslations, especially in semantically distant or morphologically rich languages (Wei et al., 2023a; Brown et al., 2020b). Current evaluation metrics such as BLEU, perplexity, or BERTScore are ill-equipped to assess riddle-specific traits like structural novelty, literary device density, or cultural fit (Sellam et al., 2020a; Dufter, 2021a).
To bridge these gaps, we propose Adaptive Originality Filtering (AOF), a prompting framework that enforces semantic novelty and lexical diversity through a cosine similarity-based rejection mechanism. Unlike typical generation strategies, AOF injects external control into the decoding loop, filtering out redundant or culturally dissonant outputs to elicit more original and resonant generations. Complementing AOF, we introduce RiddleScore, a composite evaluation metric that captures four dimensions central to high-quality riddles: Novelty, Diversity, Fluency, and Semantic Alignment. RiddleScore leverages pretrained language models alongside traditional metrics, and is calibrated to reflect human intuition across languages.
We benchmark AOF-enhanced prompting in three state-of-the-art LLMs: GPT-4o, LLaMA 3.1 and DeepSeek Reasoning in four language pairs (English,Chinese, Arabic, Japanese, French). Using the BiRdQA dataset (Zhang and Wan, 2022) under consistent decoding parameters, we evaluate outputs with Self-BLEU, Distinct-2, Cross-lingual BERTScore, and human judgment. Our results show that AOF significantly outperforms standard prompting baselines across both automatic and human evaluations. Notably, in Japanese, AOF-enhanced GPT-4o achieves a Self-BLEU of 0.177 and a Distinct-2 of 0.915, indicating reduced redundancy and heightened linguistic variety.
To structure our contributions more rigorously, we center our study around the following research questions:
  • RQ1: Can rejection-based prompting (AOF) increase semantic novelty and lexical diversity across typologically diverse languages?
  • RQ2: Does the proposed composite metric, RiddleScore, correlate with human judgments better than uniform-weighted baselines?
  • RQ3: How do pretrained versus fine-tuned LLMs respond to AOF in multilingual riddle generation?
We address RQ1 by showing that AOF with a cosine threshold of $\theta=0.75$ significantly improves novelty and diversity across languages; in Japanese, it reduces Self-BLEU to 0.177 (–63.4%) and raises Distinct-2 to 0.915. For RQ2, RiddleScore aligns strongly with human judgments (Spearman $\rho=0.83$), outperforming uniform baselines. For RQ3, we find that fine-tuned models benefit more from AOF than pretrained ones—achieving greater improvements in originality, fluency, and cultural fit. Chinese shows the most pronounced gains, with RiddleScore increasing by 48.3% (0.453 → 0.728) and human ratings rising from 3.91 to 4.50.
@@ 2 Related Work
Multilingual and Cultural NLP
Most work on riddles has focused on comprehension or solving rather than generation. Recent shared tasks such as SemEval-2024 Task 9 (Heavey et al., 2024) benchmark multilingual riddle solving with diverse unsupervised systems. RIScore (Panagiotopoulos et al., 2024) enhances contextual reasoning via in-context augmentation but does not explore generative capabilities. BiRdQA (Zhang and Wan, 2022) provides a multilingual benchmark but focuses on multiple-choice comprehension. In Chinese NLP, Xu et al. (Xu et al., 2022) incorporated cultural embeddings to improve riddle comprehension, while Tan et al. (Tan et al., 2016) explored classical Chinese radical riddles. Megatron-Turing NLG (Smith et al., 2022) includes riddles among its evaluation tasks but lacks task-specific generation. Figurative generalization remains difficult for multilingual LMs (Liu et al., 2022a), as metaphor and symbolism often fall outside pretrained representations (Dufter, 2021b). Sentence-level alignment models such as LASER (Chen and Avgustinova, 2021), XLM-R (Conneau et al., 2019), and MUSE (Lample and Conneau, 2019) improve transfer but collapse under poetic or rhetorical pressure. Our method explicitly addresses cultural fluency through semantic rejection and literary device filtering, ensuring metaphorical and idiomatic depth across languages.
Creative and Figurative Language Generation
Creative NLP tasks—such as joke generation (Petrović and Matthews, 2013), metaphor synthesis (Chakrabarty et al., 2021), and story writing (Fan et al., 2018)—highlight the tension between novelty and fluency. Studies like GENIE (Tambwekar et al., 2019) and related prompting approaches (Zhang et al., 2020a) introduce generation frameworks for idea diversity, but often lack semantic constraints. Cross-lingual creativity remains underexplored: transformer-based models (Weller and Seppi, 2019) have begun to address humor generation, yet cultural adaptation remains limited. In Chinese, visual-pun riddles require multimodal cues (Zhou and Bisk, 2022), while poetic style transfer systems like Hafez (Ghazvininejad et al., 2017) aim to generate stylized literary output. Tan et al. (Tan et al., 2016) model riddle form in character-based composition. These works suggest the need for structured prompts or heuristics to scaffold creative reasoning. Our work differs by combining cultural-device filtering with a retry loop to enforce lexical and rhetorical novelty without additional supervision.
Prompting Strategies and Constraint-Based Generation
Standard prompting methods such as few-shot and chain-of-thought (CoT) improve reasoning but tend to replicate memorized patterns (Brown et al., 2020a; Wei et al., 2023b). Recent methods like Self-Refine (Madaan et al., 2023), Reflexion (Krishna et al., 2023), and Tree-of-Thought (Yao et al., 2023) explore iterative improvement, while Auto-CoT (Zhang et al., 2022) and Selective CoT (Li et al., 2023) adapt prompt selection. Constraint-driven frameworks such as COLD decoding (Mou et al., 2022), EditCoT (Wang et al., 2024), Crescendo (Zhou et al., 2022), and Sketch-of-Thought (Aytes et al., 2025) offer structure-guided generation, but do not explicitly enforce cultural or semantic novelty. Creativity-centric methods such as SCILL (Dou et al., 2022) and CS4 (Atmakuru et al., 2024) demonstrate structure helps, but often lack filtering loops. Our Adaptive Originality Filtering framework unifies these threads by integrating rejection sampling, metaphor constraints, and interlingual filters into a single prompting loop.
Evaluation of Multilingual Generation
While BLEU and BERTScore are widely used, they poorly reflect originality or cultural fit (Dang et al., 2022; Schmidtová and Wu, 2024). BLEURT (Sellam et al., 2020b) and COMET (Rei et al., 2020) improve robustness, but do not capture rhetorical or misdirectional quality. HUME (van der Lee et al., 2021) enables human-aligned evaluation but is domain-limited. Recent surveys (van der Lee et al., 2019; Cahill et al., 2009) highlight gaps in evaluating creative NLP. Multilingual creativity requires more than fluency—fluency is necessary but not sufficient. RiddleScore, our proposed metric, captures novelty (via semantic distance), lexical diversity, fluency, and answer coherence in a single interpretable score. It extends earlier work on figurative evaluation (Shutova, 2013; Falkum, 2009) and is explicitly validated by structured human annotation across language pairs.
@@ 3 Methodology
3.1 Adaptive Originality Filtering (AOF)
To overcome shortcomings of classical prompting techniques such as Chain‑of‑Thought and Few‑Shot, which tend to copy riddles from pretraining data (Zhang and Wan, 2022), we present Adaptive Originality Filtering (AOF), a prompting technique boosting novelty, lexical richness, and cultural adherence in riddle construction.
AOF combines three core mechanisms: (1) semantic similarity filtering, (2) rejection sampling, and (3) prompt‑level constraints. For semantic filtering, a candidate riddle is encoded using MiniLM embeddings and matched to a reference set using cosine similarity. Extending from existing research where $0.75$ serves as the inflection point where topical drift becomes primarily influenced by semantic novelty (Li et al., 2024; Lee, 2025), our novelty cutoff is set to be Candidates exceeding this threshold are rejected (Appendix M.1); the full rejection‑sampling loop is given in Appendix M.2, and the prompt skeleton in Appendix M.
We verified a threshold-sensitivity study (Table 29, Appendix) that validates $\theta=0.75$ as minimizing Self-BLEU and maximizing Distinct-2, with lower thresholds that allow template bleedthrough and higher thresholds that increase the failure rate by 14 %. Figure 3 shows a visualization of the rejection-sampling loop,
Figure 2: AOF rejection-sampling loop. Each candidate is generated, compared to reference riddles, and either accepted, rejected, or retried up to $k$ attempts.
Figure 3: RiddleScore components and weights ($\alpha{=}0.30,\beta{=}0.20,\gamma{=}0.30,\delta{=}0.20$).
3.2 RiddleScore Metric
To evaluate multilingual riddle quality we introduce RiddleScore, a composite metric that captures four dimensions—Novelty, Diversity, Fluency, and Semantic Alignment. Formal definitions are in Appendix O, which also justifies the choice of the back-end models (MiniLM, Distinct-2, GPT-2.5 perplexity, and BERTScore) in a dedicated “Model Choice” paragraph.
Each component is computed as follows:
  • Novelty: cosine distance from BiRdQA riddles (MiniLM).
  • Diversity: Distinct-2 bigram ratio (Li et al., 2016).
  • Fluency: inverse perplexity under a frozen GPT-2.5 (Radford et al., 2019).
  • Semantic Alignment: BERTScore against the riddle’s answer (Zhang et al., 2020b).
The final score is a weighted sum
$\begin{split}\text{RiddleScore}=\alpha\,\text{Novelty}+\beta\,\text{Diversity}+\gamma\,\text{Fluency}+\delta\,\text{Alignment}.\end{split}$ (1)
with $\alpha\!=\!0.30$, $\beta\!=\!0.20$, $\gamma\!=\!0.30$, and $\delta\!=\!0.20$. The weights were searched by grid on a 120 sample dev set to maximize Spearman $\rho$ with 5-point human scores (Table 30); the selected setting raises $\rho$ from 0.71 (uniform) to 0.83. In addition, Appendix O Figure 9 shows how alternative weightings affect correlation with human scores. This mirrors the weight-tuning strategies of MetaMetrics (Winata et al., 2024) and HarmonicEval (Ohi et al., 2024). Figure 3 diagrams how the four components and their weights combine into riddlescore.
@@ 3.3 Experimental Setup
We test three LLMs—GPT-4o, LLaMA 3.1, and DeepSeek Reasoning—under five prompting strategies: Zero-Shot, Few-Shot, Chain-of-Thought, Adversarial (Wallace et al., 2019; Ribeiro et al., 2018), and AOF. All models are decoded with temperature $0.7$, the default in most production chat systems and evaluation suites (e.g., SORRY-Bench) and shown to balance diversity and factuality in decoding studies (Xie et al., 2025; Lu et al., 2024).
Prompts are evaluated in five languages using the BiRdQA corpus of 15 k bilingual riddles (Zhang and Wan, 2022); exact prompt templates appear in Table 25. BiRdQA is uniquely suited for this evaluation, as it captures figurative reasoning, symbolic abstraction, and cultural idiomaticity, traits essential to assess cross-lingual creativity and semantic alignment in generative models(Liu et al., 2022a; Kabra, 2023). BiRdQA has been increasingly adopted in multilingual studies as a benchmark to evaluate figurative abstraction and cross-lingual generalization(Giadikiaroglou et al., 2024; Huang et al., 2025). Our Evaluation metrics include Self-BLEU (repetition), Distinct-2 (diversity), cross-lingual BERTScore (alignment), and the composite RiddleScore. Syntactic validity is verified using spaCy and Stanza. Full metric definitions, details of datasets, and other experimental materials are included in Appendix N.
4 Fine-Tuning of the GPT-4o Model
@@
This fine-tuning was to refine solving and generating riddles in diverse languages by GPT-4o-2024-08-06. The riddles involve something beyond matching on a page—they require comprehension of metaphors, logical paradox, and novel misdirection. Our goal was to not only refine accuracy of answers but to also instill structural reasoning ability.
@@
We posed the problem as a supervised multiclass classification task with the BiRdQA dataset. The riddles were given as multiple choices, and cross-entropy loss was used to fine-tune the model. The reader can find complete details regarding dataset preprocessing, training procedure, and expanding the training set, respectively, in Appendix L.
@@
Riddles were presented as four-choice multiple-choice questions with an eye to obtaining fine-grained discrimination between plausibly believable distractors. This format affected the inference strategy and generalizability of the model. The analysis of framing effects can be seen in Appendix L.5.
@@
We tested five prompting methods on the fine-tuned model: Zero-Shot, Few-Shot, Chain-of-Thought (CoT), Adversarial, and Adaptive Originality Filtering (AOF). These correspond to the pre-trained experiments. See full prompt templates by reading Table 25.
@@
We compared our fine-tuned GPT-4o with several pre-trained baselines: GPT-4o (pre-trained), LLaMA 3.1, DeepSeek R1 with same evaluation metrics and prompts. Detailed results and discussion of methods are found in Appendix N.
Language Prompting Method Score (/5)
English AOF (Ours) 4.50
Few-Shot 3.20
Zero-Shot 3.15
Chain-of-Thought 3.85
Adversarial 4.20
Chinese AOF (Ours) 4.50
Few-Shot 3.25
Zero-Shot 3.50
Chain-of-Thought 4.00
Adversarial 3.80
Japanese AOF (Ours) 3.43
Few-Shot 3.36
Zero-Shot 3.50
Chain-of-Thought 3.57
Adversarial 3.64
French AOF (Ours) 4.44
Few-Shot 3.78
Zero-Shot 4.00
Chain-of-Thought 4.33
Adversarial 3.83
Arabic AOF (Ours) 4.92
Few-Shot 4.08
Zero-Shot 3.72
Chain-of-Thought 4.40
Adversarial 4.30
Table 1: Average human evaluation scores (out of 5) for the fine-tuned GPT-4o across languages. Best per language in bold.
Language Prompting Method Score (/5)
English AOF (Ours) 3.85
Few-Shot 2.75
Zero-Shot 2.50
Chain-of-Thought 3.52
Adversarial 3.60
Chinese AOF (Ours) 3.91
Few-Shot 2.63
Zero-Shot 2.75
Chain-of-Thought 3.45
Adversarial 3.78
Japanese AOF (Ours) 3.36
Few-Shot 2.86
Zero-Shot 2.79
Chain-of-Thought 2.93
Adversarial 3.29
French AOF (Ours) 4.50
Few-Shot 3.85
Zero-Shot 3.77
Chain-of-Thought 3.55
Adversarial 4.00
Arabic AOF (Ours) 4.92
Few-Shot 4.20
Zero-Shot 2.71
Chain-of-Thought 4.40
Adversarial 4.25
Table 2: Average human evaluation scores (out of 5) for pretrained models.
Language Pair Prompting Method GPT-4o LLaMA 3.1 DeepSeek R1 English–Arabic AOF (Ours) 0.373 0.378 0.400 Zero-Shot 0.352 0.382 0.400 Few-Shot 0.338 0.366 0.341 Adversarial 0.296 0.292 0.305 English–Chinese AOF (Ours) 0.434 0.330 0.453 Zero-Shot 0.250 0.136 0.255 Few-Shot 0.253 0.263 0.257 Chain-of-Thought 0.247 0.246 0.239 Adversarial 0.247 0.253 0.280 English–Japanese AOF (Ours) 0.367 0.341 0.379 Zero-Shot 0.351 0.363 0.323 Few-Shot 0.346 0.353 0.324 Chain-of-Thought 0.302 0.490 0.273 Adversarial 0.338 0.361 0.336 English–French AOF (Ours) 0.373 0.352 0.354 Zero-Shot 0.410 0.423 0.428 Few-Shot 0.330 0.327 0.329 Chain-of-Thought 0.236 0.350 0.241 Adversarial 0.242 0.251 0.234
Table 3: RiddleScore performance across language pairs and pretrained models.
Lang. Pair Prompting Method RiddleScore
Eng–Arabic AOF (Ours) 0.586
Few-Shot 0.364
Zero-Shot 0.315
Chain-of-Thought 0.313
Adversarial 0.341
Eng–Chinese AOF (Ours) 0.728
Few-Shot 0.355
Zero-Shot 0.350
Chain-of-Thought 0.312
Adversarial 0.348
Eng–Japanese AOF (Ours) 0.475
Few-Shot 0.334
Zero-Shot 0.300
Chain-of-Thought 0.307
Adversarial 0.331
Eng–French AOF (Ours) 0.352
Few-Shot 0.350
Zero-Shot 0.468
Chain-of-Thought 0.347
Adversarial 0.328
Table 4: Fine-tuned GPT-4o RiddleScore across language pairs.
5 Human Evaluation
To capture riddle qualities not fully represented by automatic metrics, we performed human evaluations on four axes: Fluency, Novelty, Cultural Fit, and Answerability. Native or proficient speakers rated the riddle-answer pairs on a 1- to 5-likert scale using standardized rubrics, with hidden model labels to reduce bias (Appendix P).
5.1 Results
In both pre-trained and fine-tuned models, AOF prompting achieved the highest average scores in all languages, reaching 4.92 in Arabic and 4.50 in English, Chinese and French (Tables 2 and 2). These scores substantially exceed those of the zero-shot, few-shot, and chain-of-thought prompting, demonstrating the superiority of the AOF in producing culturally grounded and semantically coherent riddles. Annotators frequently highlighted AOF’s “poetic language, cultural anchoring, and structural coherence” as reasons for higher ratings. For example, the French riddle “Dans le jardin des mots, je suis une abeille, bourdonnant entre les lettres, mais je ne pique jamais. Que suis-je?” (“In the garden of words, I am a bee, buzzing between the letters, but I never sting. What am I?”) was rated highly for its metaphorical depth and native-like phrasing, reflecting AOF’s ability to balance creativity with solvability.
Human ratings align with RiddleScore trends: languages with the highest RiddleScore under AOF—0.586 Arabic, 0.728 Chinese, 0.475 Japanese, 0.468 French, 0.586 English (Table 4)—also show the largest human-rated gains. This convergence validates RiddleScore as a reliable proxy for human perception of creativity, fluency, and cultural fit. Together, confirming AOF prompting consistently outperforms other methods.